From nch@roe.ac.uk Fri Oct 24 13:47:12 2003 Date: Fri, 24 Oct 2003 12:34:53 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Harvey MacGillivray , Ian Bond , Mike Read , Nigel Hambly , Bob Mann Cc: Andrew Lawrence , Peredur Williams , Andy Adamson , Jim Emerson , Mike Irwin Subject: WFAU WSA weekly meeting minutes: 24 October 2003 Minutes of WFCAM Science Archive meeting: 24th October 2003 ----------------------------------------------------------- ----------------------------------------------------------- Present: IAB, PMW, NCH, MAR Apologies: ETWS, JMS, AL, RGM, HMG, JPE Actions discharged: ------------------- ACTION: NCH to try to start the WSA curation populating scripts before the next meeting. (started; development scripts being checked into CVS). ACTION: NCH to put an ADASS visit report on the TWiki. (see http://apache.roe.ac.uk/twiki/bin/view/WFAU/ADASS2003Report) (thanks to all concerned). Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to get HTM2 from JHU. (Gyorgy Fekete emailed; no response yet) ACTION: ALL to review risk register and think of any new internal or external risks that should be documented in the risk register. (this one is now important again in the light of recent AGOC meetings and request from JPE to revise the register. See Project Management below). ACTION: RGM to progress design of a web cookbook based around the 20 Queries. (started) (still continues: 2/3 done) (RGM on leave from 10/10/03) ACTION: NCH, ETWS to create schema-driven tags for ingest procedure, and to code up proto-curation populating scripts for the WSA for test purposes. (schema tags to drive ingest are being implemented in revised schemas in the CVS. Proto-curation populating scripts have yet to be started) (NCH still hasn't started the curation scripts - see S/W below) (NCH now started; will need a day or so to complete). ACTION: JPE to design and set up centralised VDFS web pages. (nearly complete; see Project Management below). ACTION: NCH & PMW to meet Mon am to revisit the project plan and risk register in the light of recent AGOC advice and the WFCAM slips. (discharged, but production of required tables and GANTTs still being done) ACTION: NCH & ETWS to finalise USNOB binary load scripts and to send these and the schema to Alex at JHU. (scripts etc. done but not checked yet; will communicate the procedure to Alex once it has been tested). - continuing; thanks for progressing these. Actions carried forward from 17/10/03 meeting: ---------------------------------------------- ACTION: PMW to liaise with AL to make sure that Compute Support are made aware, from the highest "official" level, of our requirements, re. 100 Gbyte/day transfer from CASU, and to impress on them that the 1 Gbit/s FW should be their highest priority. - CONTINUES Specific points and new actions: -------------------------------- Project management: JPE has produced a nice set of centralised VDFS web pages at http://www-star.qmw.ac.uk/~jpe/vdfs The only remaining things to be done are a top-level business plan (NCH & PMW have both made suggestions to JPE concerning this). PMW & NCH met to discuss the forward plan; PMW has been producing forward plan tables (eg. a Q1'04 table of tasks for the monthly reporting) and is also looking at making a progress GANTT along the lines of the CASU one. The forward plan needs to be made consistent with the current ATC WFCAM schedule, and with any top-level plan illustrated at the VDFS web site. AL & JPE took part in the VEGA telecon (preliminary to next weeks Cambridge meeting). PMW is attending the meeting next week. Hardware: There have been no more disk errors on ahmose since updating the Seagate firmware; amenhotep is still being configured (there may be a dodgy disk on this sytem - JNTD is trying to track this down). In the meantime, a full SSA load is progressing on ahmose. A load was instigated on Mon AM, but this was interupted by a power supply glitch on Tues PM; the C1 trips cut off the power to C1 all night and the server UPSs could not cope with this (obviously). NCH pointed out that we need to investigate the functioning of the C1 trips, and also the UPS configuration, to make the servers more robust. ACTION: NCH to speak with JNTD concerning the C1 power trip switches and the WSA server UPS configuration. Software: JPE forwarded information concerning "montage" from IPAC (see also NCH's ADASS note). IAB has downloaded this and had a quick look; at first sight it seems rather difficult to use and perhaps won't handle some aspects of IR data (eg. confidence arrays). IAB reported attending the Condor seminar at NeSC, and suggested it looks promising for heavy CPU applications in the WSA - eg. spreading Jjobs over the file nodes of the mass storage system. NCH asked IAB to put a small report on the TWiki with appropriate web links. ACTION: IAB to put a small report on the NeSC Condor seminar on the TWiki. Networking: NCH & ETWS discussed implementation of CU1 on djoser, and testing the final transfer method from CASE; ETWS will progress this when he returns from leave. ACTION: ETWS to contact Guy Rixon concerning a new GridFTP test. SSA: A full load of the SSA is currently underway on ahmose; once amenhotep has stabilised, NCH will load the 10% for the indexing experiments to continue. Miscellaneous: NCH reported that it looks like the Astrowise visit will be 3/4 Dec. Currently the WFAU attendees will be NCH and PMW, but PMW suggested if anyone else would like to go in his place, that would be fine. IAB expressed interest and suggested he would consider this over the coming week. ACTION: NCH to type up and circulate these minutes. DONM: 10am, Friday 31st October, plate library.